OpenStudy (anonymous):

show that the area of a regular n-gon(polygon) is the same as that of a circle as \(n\to \infty\)

ganeshie8 (ganeshie8):

to prove : \(\large \lim \limits_{n \to \infty}~\dfrac{1}{2} n r^2 \sin(\frac{2\pi}{n}) = \pi r^2\)

OpenStudy (anonymous):

yes :) seems a bit obvious from here

ganeshie8 (ganeshie8):

yes ! sinx/x limit leads to pir^2 i hope should be easy... :)

OpenStudy (anonymous):

ohh yeah,very easy,and the next question goes>>> find the ratio of A:n if each side has length \(p\) and distance from center to corner is \(r\)

ganeshie8 (ganeshie8):

we want to find below Area ------ n is it ?

OpenStudy (anonymous):

yes ,yes,but we no longer focus on the situation n to infty

ganeshie8 (ganeshie8):

just divide n, we're done right ?

ganeshie8 (ganeshie8):

A/n = \(\large \dfrac{1}{2} r^2 \sin(\frac{2\pi}{n}) \) ?

OpenStudy (anonymous):

yes \[\frac{A}{n}=\frac{1}{2}\sin \frac {2\pi}{n}r^2\] is not enough since we still have n as a variable inside sin

OpenStudy (anonymous):

we want it expressed in terms of \[r,p\]

ganeshie8 (ganeshie8):

ohkk got u

ganeshie8 (ganeshie8):